OpenStudy (anonymous):

3/4 x something = 12

OpenStudy (anonymous):

16? Multiply by 4 3x = 48 Divide by 3 x = 16 ??

OpenStudy (anonymous):

\[\frac{3}{4}\times \text{something}=12\]multiply both sides by 4\[3\times \text{something}=48\]divide both sides by 3\[\text{something}=16\]